OpenStudy (anonymous):

Neena has 54 smiley-face stickers 36 glitter stickers and 81 heart stickers. She wants to divide them evenly among her friends. Find the greatest number tat Neena can use to divide the stickers evenly.

OpenStudy (triciaal):

She wants to divide them evenly among her friends a) She wants the same number of each type of sticker or b) how many friends does she have to give the same amount of stickers to each friend. I am thinking you mean a) but want to verify 54 = 2*27 36 = 1*27 with 9 left 81 = 3*27 27 each kind

OpenStudy (triciaal):

54 = 6*9 36 = 4*9 81 =9*9

OpenStudy (anonymous):

It dose not matter what kind and it dose not tell me how many friends.

OpenStudy (triciaal):

so go with 9 the greatest common factor of 54,36 and 81

OpenStudy (anonymous):

Thanks